# Carve3D: Improving Multi-view Reconstruction Consistency for Diffusion Models with RL Finetuning

### 21 Dec 2023 v1

@@ Abstract
Recent advancements in the text-to-3D task leverage finetuned text-to-image diffusion models to generate multi-view images, followed by NeRF reconstruction. Yet, existing supervised finetuned (SFT) diffusion models still suffer from multi-view inconsistency and the resulting NeRF artifacts. Although training longer with SFT improves consistency, it also causes distribution shift, which reduces diversity and realistic details. We argue that the SFT of multi-view diffusion models resembles the instruction finetuning stage of the LLM alignment pipeline and can benefit from RL finetuning (RLFT) methods. Essentially, RLFT methods optimize models beyond their SFT data distribution by using their own outputs, effectively mitigating distribution shift. To this end, we introduce Carve3D, a RLFT method coupled with the Multi-view Reconstruction Consistency (MRC) metric, to improve the consistency of multi-view diffusion models. To compute MRC on a set of multi-view images, we compare them with their corresponding renderings of the reconstructed NeRF at the same viewpoints. We validate the robustness of MRC with extensive experiments conducted under controlled inconsistency levels. We enhance the base RLFT algorithm to stabilize the training process, reduce distribution shift, and identify scaling laws. Through qualitative and quantitative experiments, along with a user study, we demonstrate Carve3D’s improved multi-view consistency, the resulting superior NeRF reconstruction quality, and minimal distribution shift compared to longer SFT. Project webpage: https://desaixie.github.io/carve-3d.
Figure 1: Carve3D steadily improves the 3D consistency of the multi-view diffusion model and the resulting quality of the NeRF and the mesh, without sacrificing its diversity or realism. Here we show results of the finetuned multi-view diffusion model over three epochs on three testing prompts (three blocks separated by dotted line), including the generated multi-view images (top), the reconstructed NeRF and extracted mesh (bottom) and the prompt (middle). The inconsistencies in the multi-view images, e.g. the facing direction of the shopping cart, the position of the octopus arms, and the position of the pencils, lead to artifacts in the NeRF and the mesh, highlighted by the red boxes.
@@ 1 Introduction
In recent times, notable advancements have been made in the text-to-3D domain, driven by lifting images generated by 2D diffusion models [42, 40, 51, 9, 55, 48] to 3D. Numerous methods [46, 24, 28, 58] have demonstrated that a set of multi-view images is adequate for generating diverse and detailed 3D models, effectively mitigating the Janus (two-face) problem. Thus, ensuring the 3D consistency across these multi-view images is crucial for 3D generation, as inconsistencies can inevitably introduce artifacts, such as broken geometries, blurring, or floaters, in the NeRF reconstruction. However, the lack of an established multi-view consistency metric has led researchers to rely on qualitative inspections, which are both inefficient and unreliable, highlighting the ongoing issues with inconsistency in current methods.
@@
Existing multi-view diffusion models [46, 24, 28, 58, 27] primarily utilize supervised finetuning (SFT) with datasets derived from rendering perfectly consistent multi-view images of 3D assets [12, 13]. While SFT can achieve some degree of multi-view consistency, it presents a dilemma: prolonged SFT enhances this consistency but also induces a distribution shift that diminishes diversity and realism of the results [24]. Such dilemma has been observed in the research of large language models (LLMs). While SFT changes the output distribution of pre-trained LLMs from text completion to answering instructions, such distribution shift to the instruction dataset also introduces hallucination [44], preventing longer SFT. InstructGPT [36], the paper behind ChatGPT 3.5 [35], introduces RL finetuning (RLFT) to further align the SFT model without causing additional distribution shift. Drawing an analogy between instruction-finetuned LLMs and multi-view diffusion models, RLFT emerges as an essential step following the SFT stage. By adopting RLFT, we aim to enhance the consistency of multi-view diffusion models without introducing the biases from a SFT dataset (Figure 1).
@@
We introduce Carve3D, a RLFT framework paired with our Multi-view Reconstruction Consistency (MRC) metric, to improve the consistency of multi-view diffusion models (Figure 2). MRC compares the output multi-view images from a diffusion model, which serve as inputs for NeRF reconstruction, with images rendered from the NeRF at identical camera viewpoints. We use the sparse-view Large Reconstruction Model (LRM) [17, 24] to achieve fast, feed-forward NeRF reconstruction from a few multi-view images. To quantify image similarity, we adopt LPIPS [56] as it is more effective and robust for MRC. We further normalize LPIPS with respect to the bounding boxes of foreground objects to prevent trivial reward hacking through size reduction of the foreground object. To validate the reliability of MRC, we conduct extensive experiments with controlled inconsistency levels; starting from a set of perfectly consistent multi-view images rendered from a 3D asset [12], we manually introduce distortion to one of the views to create inconsistency. Our MRC metric provides robust evaluation of consistency of multi-view images, offers a valuable tool for assessing current multi-view generation methods and guiding future developments in the field.
@@
With MRC, we employ RLFT for multi-view diffusion models. In the RLFT process, we use a set of curated, creative text prompts to repeatedly generate diverse multi-view images with random initial noises and use their MRC reward to update the diffusion model (Figure 2). Such diversity- and quality-preserving finetuning cannot be achieved with SFT, as it is infeasibly expensive to create a dataset of diverse ground-truth multi-view images for these prompts. We make the following improvements to the RLFT algorithm [5]. In addressing the common issue of training instability in RL, we opt for a purely on-policy policy gradient algorithm [53], diverging from the widely adopted, partially on-policy PPO [45] algorithm. We incorporate KL divergence regularization [15, 36] to maintain proximity to the base model and prevent distribution shift. Moreover, we scale up the amount of compute to achieve optimal rewards by applying the scaling laws for diffusion model RLFT, identified from extensive experiments – a topic that has not yet been extensively covered in existing studies [5, 15].
@@
Through quantitative and qualitative experiments, as well as human evaluation, we demonstrate that Carve3D: (1) achieves improved multi-view consistency and NeRF reconstruction quality over the base multi-view diffusion model, Instant3D-10K [24], as well as Instant3D-20K and Instant3D-100K, which utilize more SFT steps, and (2) maintains similar prompt alignment, diversity, and realistic details from the base Instant3D-10k, preventing the degradation in Instant3D-20k and Instant3D-100k. We extend our consistency evaluation to additional multi-view diffusion models using MRC, revealing the universal presence of multi-view inconsistency when relying solely on SFT. Our work is the first application of RLFT to the text-to-3D task, especially with diffusion models at the SDXL [39] scale using a 2.6B-parameter UNet. We hope this work will bolster the research on RLFT for alignment in the computer vision community.
@@
Figure 2: Overview of Carve3D. Given a prompt sampled from the curated prompt set, we run the denoising process to obtain the final denoised image, which contains four multi-view images tiled in a 2-by-2 grid. MRC reward is computed by comparing (a) the generated multi-view images with (c) the corresponding multi-view images rendered at the same camera viewpoints from (b) the reconstructed NeRF. Then, we train the model with policy gradient loss function, where the loss is derived from the reward and log probabilities of the model’s predicted noise, accumulated across all denoising timesteps. Using only a set of prompts, this RLFT process finetunes the diffusion model with its own outputs, without relying on ground truth multi-view images.
@@ 2 Related Works
NeRF is a neural representation of 3D assets [31, 8, 33]. It infers the direction-dependent radiance at an arbitrary volumetric position with neural models. Many text-to-3D methods rely on NeRF to produce 3D objects.
@@
While text-to-image diffusion models are trained on 5 billion data [43], the largest public 3D dataset only contains 10 million 3D assets [12, 13] with little text annotation. This gap in the diversity and quality of 3D data has restricted the quality of current 3D diffusion models and their ability in handling complex prompts [34, 19]. To circumvent this limitation, another line of work focuses on lifting 2D images to 3D, thus leveraging the remarkable semantic understanding and high-quality generation capabilities of 2D diffusion models [42, 39]. These methods [40, 51, 9, 55, 48] typically employ 2D diffusion models to provide supervision at the novel views for optimizing 3D objects represented as NeRF or by 3D Gaussian Splatting [21]. Building on this concept, multiple works [27, 26, 46, 24, 28, 58] have proposed generating multi-view images using a finetuned 2D diffusion model, providing a more comprehensive visual prior and preventing the multi-face (Janus) problem. However, as the finetuning datasets of multi-view images are rendered from the same 3D dataset [12, 13], the limited quality and diversity remains a challenge, preventing running SFT to convergence [24]. By adopting RLFT, we do not depend on ground truth multi-view images and thus optimize the model beyond the distribution of their SFT dataset.
@@
A key challenge in utilizing multi-view images is achieving 3D consistency, ensuring that the geometry and appearance of an object is uniform across different views. While numerous methods have attained notable multi-view consistency by supervised finetuning 2D diffusion models [46, 24, 28, 58, 27], their evaluation has been empirical, lacking explicit metrics. An approach known as 3D consistency scoring [52] measures the consistency of output views by optimizing a NeRF trained on these views. However, this method requires dense input view sampling for cross-validation, making it unsuitable for evaluating sparse views. To overcome this limitation, we propose the MRC metric. MRC evaluates multi-view consistency by comparing input images with renderings of the generated 3D objects from corresponding views. We employ LPIPS, a perceptual image distance metric [56], to quantify the image differences. Additionally, we validate the reliability of MRC by conducting extensive experiments under controlled inconsistency levels. The robustness of MRC allows us to conduct RLFT on multi-view diffusion models, which significantly enhances their multi-view consistency.
@@
3D models can be derived from either single or multi-view images by optimizing the SDS distillation loss [40, 51]. However, the optimization process is notably time-consuming, requiring multiple hours to generate a single 3D asset. In contrast, LRM [17], trained on the extensive 3D dataset Objaverse [12], can efficiently reconstruct NeRF models from a single image in a feed-forward manner. In this work, we focus exclusively on text-to-3D using feed-forward sparse-view NeRF reconstruction, specifically employing sparse-view LRM [24]. This choice is driven by its significantly faster performance compared to SDS-based optimization methods and its superior quality relative to feed-forward text-to-3D diffusion models [34, 19]. We choose Instant3D [24] as our base multi-view diffusion model, owing to its light-weight SFT that preserves the strong semantic understanding and high-quality image generation capabilities of SDXL [39], similar to the instruction finetuning stage in InstructGPT [36].
@@ 2.2 RLFT of LLMs and Diffusion Models
RL has been widely used to finetune large pre-trained models in NLP [36, 3, 22, 2] and CV [5, 38, 15, 41, 11, 57], due to its advantage over SFT. SFT directly fits a model to the distribution of the SFT dataset containing inputs and ground-truth target data, which unavoidably causes some degree of distribution shift [44]. On the contrary, based on an objective function and a dataset containing only inputs, RLFT optimizes a model beyond the limitation of a SFT dataset by using its own outputs and effectively mitigates distribution shift [7].
@@
LLMs like GPT-3 [6] are pre-trained on the next-word prediction task on an internet-scale corpus. While the autoregressive pre-training is a powerful self-supervised objective that allows LLMs to extract substantial knowledge from the internet-scale unlabeled dataset, pre-trained LLMs can only perform the corresponding text completion task. The pre-training lacks an objective that allows LLMs to respond to text prompts. In InstructGPT [36], the paper behind ChatGPT 3.5, a two-stage finetuning solution is proposed to align GPT-3 to answer instructions according to human preferences. In the first stage, InstructGPT employs SFT with a small dataset of hand-crafted prompt-answer pairs. While SFT changes the model’s output distribution from text completion to answering instructions, it also introduces hallucination [44]. This is because the output distribution drifts too much towards the instruction-following dataset, and the model tries to imitate the behavior in the data and always provide plausible answers even when the model is uncertain about the answer [44]. To address this issue, InstructGPT opts for a light-weight SFT stage and relies on RLFT in the second stage, using a human-preference reward model. This approach provides general alignment to human values and causes minimal hallucination [44], because RLFT does not rely on a potentially biased dataset containing fixed ground-truth answers, but instead learns the general concept of human-preference through the reward model. The success of InstructGPT [36] and its analogy to the distribution shift problem in multi-view SFT [24] motivate us to pursue RLFT for 2D diffusion models.
@@
Figure 3: Qualitative correlation between MRC and inconsistency with increasing intensity, introduced by inpainting with increasing mask sizes. Left: the four ground truth views. Right: the 4th view is inpainted with increasing area sizes, i.e. 0$\times$0, 50$\times$50 and 110$\times$110 pixels. The top row is the image after inpainting and the bottom row is the image rendered from the NeRF reconstructed with the top inpainted 4th view and the other 3 original views. We mark the inpainting area with blue and red boxes. Since the lion’s right paw in the inpainted 4th views look different from the other three original views, its geometry is broken in the NeRF rendered views. This difference is captured in MRC’s image dissimilarity metric.
Figure 4: Quantitative correlation between MRC and inconsistency with increasing intensity, for the object shown in Figure 3. As inconsistency intensity rises, MRC also monotonically increases.

### 09 Apr 2024 v2

@@ Abstract
Multi-view diffusion models, obtained by applying Supervised Finetuning (SFT) to text-to-image diffusion models, have driven recent breakthroughs in text-to-3D research. However, due to the limited size and quality of existing 3D datasets, they still suffer from multi-view inconsistencies and Neural Radiance Field (NeRF) reconstruction artifacts. We argue that multi-view diffusion models can benefit from further Reinforcement Learning Finetuning (RLFT), which allows models to learn from the data generated by themselves and improve beyond their dataset limitations during SFT. To this end, we introduce Carve3D, an improved RLFT algorithm coupled with a novel Multi-view Reconstruction Consistency (MRC) metric, to enhance the consistency of multi-view diffusion models. To measure the MRC metric on a set of multi-view images, we compare them with their corresponding NeRF renderings at the same camera viewpoints. The resulting model, which we denote as Carve3DM, demonstrates superior multi-view consistency and NeRF reconstruction quality than existing models. Our results suggest that pairing SFT with Carve3D’s RLFT is essential for developing multi-view-consistent diffusion models, mirroring the standard Large Language Model (LLM) alignment pipeline. Our code, training and testing data, and video results are available at: https://desaixie.github.io/carve-3d.
Figure 1: Our Carve3D algorithm steadily improves the 3D consistency of a multi-view diffusion model and the resulting quality of the NeRF and the mesh, without sacrificing its image-prompt alignment, texture details, or realism. Here, we show 3 testing-set results (in 3 rows, numbered as 1-3, separated by dotted lines) from the finetuning process (epoch 0, 28, and 55 in 3 columns). Each row includes the generated multi-view images (denoted as MV), the reconstructed NeRF and extracted mesh (denoted as RM) and the text prompt (denoted as TP). The inconsistencies in the multi-view images, e.g. the facing direction of the shopping cart, the position of the octopus arms, and the position of the pencils, lead to artifacts in the NeRF and the mesh (highlighted in red).
†This work is done while the author is an intern at Adobe Research.
@@ 1 Introduction
Recently, significant progress has been made in generating 3D models from text prompts. Images generated by 2D diffusion models  [42, 40, 52, 9, 56, 49] can be lifted to 3D representations. Numerous methods [46, 24, 28, 59] have demonstrated that a set of multi-view images is adequate for generating diverse and detailed 3D models, effectively mitigating the multi-face (Janus) problem. Ensuring the 3D consistency across these multi-view images is crucial for 3D generation, as inconsistencies can inevitably introduce artifacts, such as broken geometries, blurring, or floaters, in the Neural Radiance Field (NeRF) reconstruction. However, the lack of an established multi-view consistency metric has led researchers to rely on qualitative inspections, which are both inefficient and unreliable.
@@
Existing multi-view diffusion models [46, 24, 28, 59, 27] primarily utilize Supervised Finetuning (SFT) with multi-view datasets derived from 3D datasets [12, 13]. While SFT can achieve some degree of multi-view consistency, it presents a dilemma: prolonged SFT enhances multi-view consistency but also induces a distribution shift towards the 3D dataset, which has limited size and quality; thus, the distribution shift diminishes diversity, texture details, and realism of the generated results [24]. Such dilemma has been observed in Large Language Model (LLM) research. While SFT can shift the output distribution of pre-trained LLMs to follow instructions, the bias from the instruction dataset also introduces hallucination [44], preventing longer SFT. InstructGPT [36], the paper behind ChatGPT 3.5 [35], introduces Reinforcement Learning finetuning (RLFT) to further align the SFT model without causing additional distribution shift. Drawing an analogy between instruction-finetuned LLMs and multi-view diffusion models, RLFT emerges as an essential step following the SFT stage. By adopting RLFT, we aim to enhance the consistency of multi-view diffusion models without introducing the bias from a SFT dataset.
We introduce Carve3D, an enhanced RLFT algorithm paired with a novel Multi-view Reconstruction Consistency (MRC) metric, to improve the consistency of multi-view diffusion models. Figs. 1 and 2 shows the capability and overview of Carve3D.
@@
Our MRC metric compares the output multi-view images from a diffusion model with images rendered from the reconstructed NeRF at identical camera viewpoints. We use the sparse-view Large Reconstruction Model (LRM) [17, 24] to achieve fast, feed-forward NeRF reconstruction from a few multi-view images. To quantify image similarity, we adopt LPIPS [57] as it is more effective and robust for MRC. We further normalize LPIPS with respect to the bounding boxes of foreground objects to prevent trivial reward hacking through size reduction of the foreground object. To validate the reliability of MRC, we conduct extensive experiments with controlled inconsistency levels; starting from a set of perfectly consistent multi-view images rendered from a 3D asset [12], we manually introduce distortion to one of the views to create inconsistency. Our MRC metric provides robust evaluation of consistency of multi-view images, offers a valuable tool for assessing current multi-view generation methods and guiding future developments in the field.
@@
With MRC, we employ RLFT for multi-view diffusion models. In the RLFT process, we use a set of curated, creative text prompts to repeatedly generate diverse multi-view images with random initial noises and use their MRC reward to update the diffusion model (Fig. 2). Such diversity- and quality-preserving finetuning cannot be achieved with SFT, as it is infeasibly expensive to create a dataset of diverse ground-truth multi-view images for these prompts. We make the following improvements to the RLFT algorithm [5]. In addressing the common training instability issue in RL, we opt for a purely on-policy policy gradient algorithm [54], diverging from the widely adopted, partially on-policy PPO [45] algorithm. We incorporate KL divergence regularization [15, 36] to maintain proximity to the base model and prevent distribution shift. Moreover, we scale up the amount of compute to achieve optimal rewards by applying the scaling laws for diffusion model RLFT, identified from extensive experiments – a topic that has not yet been extensively covered in existing studies [5, 15].
@@
By adopting our Carve3D RLFT algorithm on Instant3D-10K [24], a multi-view diffusion model supervised finetuned from SDXL [39], we obtain the Carve3D Model (Carve3DM). Through quantitative, qualitative experiments and a user study, we demonstrate that Carve3DM: (1) achieves improved multi-view consistency and NeRF reconstruction quality over Instant3D-10K, -20K, and -100K models, and (2) maintains similar prompt alignment, diversity, and realistic details as the base Instant3D-10K, preventing the degradation in Instant3D-20K and -100K. Our results indicate that pairing SFT with Carve3D’s RLFT is essential for developing multi-view consistent diffusion models. In addition, we extend our MRC evaluation to existing models, revealing the universal presence of multi-view inconsistency when relying solely on SFT. Our work is the first application of RLFT to text-to-3D, especially on a 2.6B-parameter denoising UNet from SDXL [39]. By releasing our code, training and testing data, we hope this work will bolster the RLFT and alignment research in the computer vision community.
@@
Figure 2: Overview of Carve3D. Given a prompt sampled from our curated prompt set and a initial noisy image, we iteratively denoise the image using the UNet. The final, clean image contains four multi-view images tiled in a 2-by-2 grid. MRC reward is computed by comparing (a) the generated multi-view images with (c) the corresponding multi-view images rendered at the same camera viewpoints from (b) the reconstructed NeRF. Then, we train the model with policy gradient loss function, where the loss is derived from the reward and log probabilities of the UNet’s predictions, accumulated over all denoising timesteps. By using only a set of training text prompts, our RLFT algorithm finetunes the diffusion model by evaluating its own generated outputs, without relying on ground truth multi-view images.
@@ 2 Related Works
Neural Radiance Field (NeRF) is a neural representation of 3D assets [31, 8, 33]. It infers the direction-dependent radiance at arbitrary volumetric positions with neural models. Many text-to-3D methods rely on it to produce 3D objects.
@@
While text-to-image diffusion models are trained on 5 billion data [43], the largest public 3D dataset only contains 10 million 3D assets [12, 13] with little text annotation. This gap in the diversity and quality of 3D data has restricted the quality of current 3D diffusion models and their ability in handling complex prompts [34, 19]. To circumvent this limitation, another line of work focuses on lifting 2D images to 3D, thus leveraging the remarkable semantic understanding and high-quality generation capabilities of 2D diffusion models [42, 39]. These methods [40, 52, 9, 56, 49] typically employ 2D diffusion models to provide supervision at the novel views for optimizing 3D objects represented as NeRF or by 3D Gaussian Splatting [21]. Building on this concept, multiple works [27, 26, 46, 24, 28, 59] have proposed generating multi-view images using a finetuned 2D diffusion model, providing a more comprehensive visual prior and preventing the multi-face (Janus) problem. However, as the finetuning datasets of multi-view images are rendered from the same 3D dataset [12, 13], the limited quality and diversity remains a challenge, preventing running Supervised Finetuning to convergence [24]. By adopting Reinforcement Learning Finetuning (RLFT), we do not depend on ground truth multi-view images and thus optimize the model beyond the distribution of their SFT dataset.
@@
A key challenge in utilizing multi-view images is achieving 3D consistency. While numerous methods have attained notable multi-view consistency by supervised finetuning 2D diffusion models [46, 24, 28, 59, 27], their evaluation has been empirical, lacking explicit metrics. An approach known as 3D consistency scoring [53] measures the consistency of output views by optimizing a NeRF trained on these views. Our MRC metric improves it with sparse-view reconstruction and comparing on all input views.
@@
3D models can be derived from either single or multi-view images by optimizing the Score Distillation Sampling (SDS) loss [40, 52]. However, the optimization process is notably time-consuming, requiring multiple hours to generate a single 3D asset. In contrast, Large Reconstruction Model (LRM) [17], trained on the extensive 3D dataset Objaverse [12], can efficiently reconstruct NeRF models from a single image in a feed-forward manner. In this work, we focus exclusively on text-to-3D using feed-forward sparse-view NeRF reconstruction, specifically employing sparse-view LRM [24]. This choice is driven by its significantly faster performance compared to SDS-based optimization methods and its superior quality relative to feed-forward text-to-3D diffusion models [34, 19]. We choose Instant3D [24] as our base multi-view diffusion model, owing to its light-weight Supervised Finetuning (SFT) that preserves the strong semantic understanding and high-quality image generation capabilities of SDXL [39], similar to the instruction finetuning stage in InstructGPT [36].
@@ 2.1 RLFT of LLMs and Diffusion Models
Reinforcement Learning (RL) has been widely used to finetune large pre-trained models in Natural Language Processing [36, 3, 22, 2] and Computer Vision [5, 38, 15, 41, 11, 58], due to its advantage over Supervised Finetuning (SFT). SFT directly fits a model to the distribution of the SFT dataset containing inputs and ground-truth target data, which unavoidably causes some degree of distribution shift [44]. On the contrary, based on an objective function and a dataset containing only inputs, Reinforcement Learning Finetuning (RLFT) optimizes a model beyond the limitation of a SFT dataset by using its own outputs and effectively mitigates distribution shift [7].
@@
Large Language Models (LLMs) like GPT-3 [6] are pre-trained on the next-word prediction task on an internet-scale corpus. While the autoregressive pre-training is a powerful self-supervised objective that allows LLMs to extract substantial knowledge from the internet-scale unlabeled dataset, pre-trained LLMs can only perform the corresponding text completion task. The pre-training lacks an objective that allows LLMs to respond to text prompts. In InstructGPT [36], the paper behind ChatGPT 3.5, a two-stage finetuning solution is proposed to align GPT-3 to answer instructions according to human preferences. In the first stage, InstructGPT employs SFT with a small dataset of hand-crafted prompt-answer pairs. While SFT changes the model’s output distribution from text completion to answering instructions, it also introduces hallucination [44]. This is because the output distribution drifts too much towards the instruction-following dataset, and the model tries to imitate the behavior in the data and always provide plausible answers even when the model is uncertain about the answer [44]. To address this issue, InstructGPT opts for a light-weight SFT stage and relies on RLFT in the second stage, using a human-preference reward model. This approach provides general alignment to human values and causes minimal hallucination [44], because RLFT does not rely on a potentially biased dataset containing fixed ground-truth answers, but instead learns the general concept of human-preference through the reward model. The success of InstructGPT [36] and its analogy to the distribution shift problem in multi-view SFT [24] motivate us to pursue RLFT for 2D diffusion models.
@@
Figure 3: Qualitative correlation between MRC and multi-view inconsistency with increasing intensity, introduced by inpainting with increasing mask sizes. Left: the four ground truth views. Right: the 4th view is inpainted with increasing area sizes, i.e. 0$\times$0, 50$\times$50 and 110$\times$110 pixels. The top row is the image after inpainting and the bottom row is the image rendered from the NeRF reconstructed with the top inpainted 4th view and the other 3 original GT views. We mark the inpainting area with blue and red boxes. Since the lion’s right paw in the inpainted 4th views look different from the other three original views, its shape is broken in the NeRF and the rendered views. This difference is captured in MRC’s image dissimilarity metric.
Figure 4: Quantitative correlation between MRC and multi-view inconsistency with increasing intensity, for the object shown in Figure 3. As inconsistency intensity rises, MRC also monotonically increases.